OpenStudy (anonymous):

Solve 3x+1 = 500 I don't know how to do this. :C

OpenStudy (anonymous):

3^(x+1)=500

OpenStudy (anonymous):

what do you mean by 3^ part??

OpenStudy (anonymous):

3x+3=500 sub 3 3x=497 divide 3 x=497/3 or 165.7